CRY+VIS Interface Meeting
Date: 2021/10/1 13:00-13:40
Participants: Takahashi, Ikeda, Yokozawa, Uchiyama, Hirata, Akutsu, Ushiba, Sato, Aso, Sugiyama, Node
- IX: Health check. Effect due to the convection was investigated. LVDT driver for the IP was replaced. Geophone signals are drifting.
- EX: Health check. Installation of ACCs will be scheduled after the health check.
- IY: H4 geophone was assembled. Yaw picomotor in BF is not working.
- EY: Cables for the BF damper were repaired.
- TM height setting: ITMY is too low.
- Commissioning in air: Health check.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
- Heater control: Al tape for the heater and bag for the adapter were implemented in IX.
Date: 2021/10/1 14:00-14:40
Participants: Takahashi, Fabian, Hirata, Sato, Ushiba, Miyoki, Aso, Terrence, Uchiyama, Sugiyama, Tanaka, Ishizaki
- Replaced the LVDT driver for IP/F0 in IX. The LVDTs were calibrated.
- Adjusted the yaw of BF damper and the tilt of IP in IY.
- Repaired cables for the BF damper in EY.
- Assembled the H4 geophone for IY. This will be used for the measurement in OMC temporally.
Investigated the yaw picomotor in IY. One screw was wedged between the keystone and the receptacle, preventing the Yaw from rotating. Screw.jpg
- Check the geophones in IY.
- SR2: we fixed cables
- The coil in IM-V3 OSEM had a larger resistance between pins 2 and 7. Replacing the pins in a male connector solved the problem.
- A coil in the F0 yaw stepper motor had a larger resistance than expected. We fixed it by crimping some pins more strongly.
- A coil in the F0 yaw stepper motor had a larger resistance than expected. A couple of weeks ago we detected the problem was with pins connecting the motor cables to an extension cable. At that time we pushed the pins together and the problem dissapear. This week we were not able to detect a problem. It seems the connection is healthy.
- Aso-san prepared a solution for the problem with the F1 primary coil. This is a discussion topic.
- We found that the flip adapter of the BF picomotors was not working, so we replaced it.
- We examined the recent change the mirror pitch due to the Sep. 19th earthquake.This is a discussion topic.
- Type-A team: 10/4: Installation of geophone to OMC, 10/4-6 Recovery of yaw picomotor in IY, 10/5 Adjustment of IP, 10/7 Replacement of LVDT driver in EX.
- Type-B team: Recovery from PR3 jump. Test of outer frame in BS.
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
- T3 Red in Y-arm had an accident.
- [Fabian] PR3 mirror pitch change due to the earthquake.
- [Hirata-san] Geometric information about PR3 payload.
- [Aso-san] Solution for the problem with the BS F0 LVDT primary coil.
(Date in Kamioka)
The week from 10/4
- Takahashi: 10/4-8
- Sato: 10/4-8
- Hirata: 10/5-8
- Yano: 10/5-8
The week from 10/11
- Takahashi: 10/12-15
- Sato: 10/12-15
- Hirata: 10/12-15
- Yano: 10/12-15
- Aso: 10/12-15